40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 26 17:50:22 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
    Reply  Wed Jan 27 18:19:45 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
       Reply  Thu Jan 28 00:29:30 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement PLL_OLG.pdfFreq_noise_comparison.pdf
          Reply  Mon Feb 1 18:11:25 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement Free_running_frequency_noise_comparison.pdf
       Reply  Fri Jan 29 14:43:47 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.pdf
          Reply  Fri Jan 29 16:55:27 2016, rana, Update, Green Locking, Innolight laser is 10 years old 
             Reply  Sat Jan 30 00:12:22 2016, gautam, Update, Green Locking, Innolight laser is 10 years old RIN_comparison.pdf
                Reply  Sat Jan 30 09:56:24 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                   Reply  Mon Feb 1 15:16:28 2016, gautam, Update, Green Locking, Innolight laser is 10 years old 
                      Reply  Mon Feb 1 15:43:18 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                         Reply  Fri Feb 5 15:02:13 2016, gautam, Update, Green Locking, X-end NE cable 
          Reply  Fri Feb 12 11:10:49 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.jpg
          Reply  Mon Mar 21 14:29:32 2016, Steve, Update, Calibration-Repair, 1W Innolight laser repair diagnoses 
             Reply  Thu Mar 24 07:56:09 2016, Steve, Update, Calibration-Repair, NO Noise Eater for 1W Innolight 
                Reply  Mon Apr 11 17:03:41 2016, Steve, Update, Calibration-Repair, 1W Innolight repair completed 
Message ID: 11967     Entry time: Mon Feb 1 15:16:28 2016     In reply to: 11964     Reply to this: 11968
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Innolight laser is 10 years old 

The Innolight laser control unit has a 25 pin D-sub connector on the rear which is meant to serve as a diagnostics aid, and the voltages at the various pins should tell us the state of various things, like the diode power monitor, laser crystal TEC error temperature, NE status etc etc. Unfortunately, I am unable to locate a manual for this laser (online or physical copy in the filing cabinets), so the only thing I have to go on is a photocopied page that Steve had obtained sometime ago from the manual for the 2W NPRO. According to that, Pin 1 is "Diode laser 1, power monitor, 1V/W". The voltage I measured (with one of the 25 pin breakout boards and a DMM) is 1.038V. I didn't see any fast fluctuations in this value either. It may be that the coefficient indicating "normal" state of operation is different for the 1W model than the 2W model, but this measurement suggests the condition of the diode is alright after all?

I also measured the voltage at Pin 12, which is described in the manual as "Noise Eater, monitor". This value was fluctuating between ~20mV and ~40mV. Toggling the NE switch on the front of the control unit between ON and OFF did not change this behaviour. The one page of the manual that we have, however, doesnt provide any illumination on how we are supposed to interpret the voltage measured at this pin...

ELOG V3.1.3-